OpenStudy (anonymous):

Input the equation of the given line in standard form. The line through (1, -4) and parallel to 2x +3y = 4.

OpenStudy (mikurout):

1st find out the slope of the given line.

OpenStudy (midhun.madhu1987):

Write the equation 2x + 3y = 4 in the form : y = mx + c and u'll get m.

OpenStudy (anonymous):

I tkink that line of equation have form 2x +3y = K If point (1, -4) lies on 2x +3y = K, K = 2*1+3*(-4) = 2-12=-10 So answer is: 2x +3y = -10 Explanation: normal vector of first line is (2,3) and paralel line have same normal vector;)